# From Sequence to Structure: Uncovering Substructure Reasoning in Transformers

### v1

@@ 3.2.2 Question prompt encoders
4 Consistency of Substructure extraction in LLMs and Transformers
As discussed in Section 3, we identify three key findings: In terms of the Transformer mechanism, (1) Transformers perform the ISF process to extract substructures simultaneously. Regarding input formulation, (2) both EL and AL can represent the adjacency matrix $A(G)$, though EL may perform slightly worse than AL due to sequence length limitations, and (3) Transformers tend to abstract substructures into a sequence of tokens in the question prompt, rather than fully capturing the underlying topological concepts. Since decoder-only Transformers are a common architecture in modern LLMs, we further evaluate whether LLMs exhibit the ISF process during substructure extraction, and whether our understanding of input formulation can explain their behavior.
@@ 4 Consistency of Substructure extraction in LLMs and Transformers
To investigate whether LLMs exhibit the ISF process, we visualize the fine-tuned LLaMA 3.1-8B-Instruct model on a triangle detection task (details in Section B.3.2), as shown in Figure 11. The results align with our findings in Section 3.1: the model often identifies the correct answer before generating it, and deeper layers better distinguish between similar answers. For example, the responses of the substructures in node ID ‘302’ and ‘304’ become more separable at layer 23 than at layer 17. We quantify this trend in Figure 11, where Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI) scores increase with depth. However, unlike Transformers trained from scratch, which only predict answers, LLMs can generate explanatory content. For instance, 23% of responses include Python code, while others describe node relationships in the graph. As a result, ARI and NMI scores slightly decrease in the ending layers.
@@
5 Implications of Substructure Extraction in Transformers
In Section 3, we highlight the ISF process as a key mechanism enabling Transformers to solve the substructure extraction task. Moreover, we observe that when text-based descriptions can be mapped to the adjacency matrix $A(G)$, Transformers can perform substructure extraction on the given graph. Building on these insights, we explore broader applications: in Section 5.1, we introduce a new method for efficiently reasoning over composite substructures and in Section 5.2, we examine how Transformers adapt to attributed graphs.
@@ 5.1 Thinking in substructures
As introduced in Section 3, Transformers apply the ISF process to extract substructures. This process runs synchronously across different patterns, with smaller ones being easier to detect. Besides, complex structures often consist of simpler components, which we refer to as decomposing substructures. With these two facts, we propose the Thinking-in-substructure (Tins) method for solving complex substructures. For example, [6] reports that reasoning language models decompose a house pattern into a triangle and a square to search for the substructures. We reformulate the answer generation part as $\mathsf{ANS_{Tins}}(\hat{G})=(\{P_{1}\},\{P_{2}\},\dots,\{P_{t}\},\mathrm{<ANS>},\mathsf{ANS}(\hat{G}))$, where $\{P_{i}\}$ is the collection of each decomposing structure and $\mathrm{<ANS>}$ is a special token to indicate that the followings are final answers. This decomposition reduces the extraction complexity from $O(n^{k})$ to $O(n^{q})$, where $q$ and $k$ are the maximum size of decomposing substructures and target substructures, respectively, with $q<k$. We show the proof in Theorem A.13 in Section A.4.
To verify the efficiency of Tins, in the experiment, we design 4 different composite substructures with their decomposition process trained with 100K samples. The experiment settings are in Section B.3.3. The results suggest that Tins can help transformers significantly improve the performance with limited training data. The overall performance can increase 10%, and in the 3 layer for diagonal structure, the performance increases about 46% percent.
@@
Substructures Directly Preds Decomposation Tins
4 layers 3 layers 4 layers 3 layers
Diagnoal [Uncaptioned image] 0.6314 0.1998    [Uncaptioned image] 0.8648 0.6606
@@ 5.2 Attributed graphs
As shown in Theorem 3.5, the only condition required for Transformers to extract substructures is $\mathcal{T}(G,G^{\prime})=1$. Therefore, if node features are uniquely assigned to ensure a distinct graph representation for each graph, Transformers can extract substructures while incorporating these features, as discussed in Theorem A.16. Taking the AL-based feature description as an example, we define the attributed graph representation as $\mathsf{AL_{f}}(G)=(v_{1}f_{1};\text{``:''};v_{1}^{1}f_{1}^{1};\cdots;v_{1}^{m_{1}}f_{1}^{m_{1}};\text{``,''};\cdots;\text{``,''};v_{n}f_{n};\text{``:''};v_{n}^{1}f_{n}^{1};\cdots;v_{n}^{m_{n}}f_{n}^{m_{n}}).$ , where $f_{i}$ is the node features.
@@ 6 Related work
Theory now probes how vanilla Transformers perform graph reasoning. Log-depth models suffice, and are necessary, for connectivity and cycles [15], while width can trade for depth [27]. ALPINE [22] shows a GPT layer embeds adjacency and reachability, validating on planning tasks; two-layer decoders trained on shortest-path learn spectral line-graph embeddings instead of Dijkstra-style rules [4]. Surveys relate attention power to Weisfeiler–Lehman bounds and over-squashing limits [12, 17]; scalable variants such as AnchorGT mitigate $O(n^{2})$ cost without losing accuracy [31]. Hierarchical distances or sparse global attention keep Transformers competitive on large or molecular graphs [8, 7, 28]. Collectively, these studies view attention heads as induced-neighbourhood selectors—exactly the mechanism ISF formalises via filtration depth.
@@ 7 Conclusion
Appendix A Appendix: Theoretical Analysis
A.1 Preliminaries
@@
Definition A.1 (Adjacency List Graph Representation).
@@
Definition A.2 (Edge List Graph Representation).
@@
Definition A.3 (Tensor Vectorization).
@@
Definition A.5 (Subgraph Isomorphism Indicator Tensor).
@@
$\mathcal{T}(G,G^{\prime})_{j_{1},j_{2},\dots,j_{k}}\begin{cases}=1,&\text{if the mapping }f:V^{\prime}\to V\text{ defined by }f(v^{\prime}_{l})=v_{j_{l}}\text{ for }l=1,\dots,k\\ &\text{satisfies both conditions:}\\ &\quad\text{(i) Injectivity: }v_{j_{1}},v_{j_{2}},\dots,v_{j_{k}}\text{ are distinct vertices in }V\\ &\quad\quad\text{(i.e., }j_{l}\neq j_{m}\text{ for all }1\leq l<m\leq k).\\ &\quad\text{(ii) Edge Preservation: For every directed edge }(v^{\prime}_{p},v^{\prime}_{q})\in E^{\prime},\\ &\quad\quad\text{the directed edge }(f(v^{\prime}_{p}),f(v^{\prime}_{q}))=(v_{j_{p}},v_{j_{q}})\text{ exists in }E.\\ \leq 0,&\text{otherwise.}\end{cases}$
@@
Throughout our theoretical analysis, we consider log-precision auto-regressive Transformer, instead of constant-precision Transformer. See [11, Appendix B] for more discussions.
A.2 Technical Lemmas
Lemma A.6 (Adjacency Matrix Extraction).
@@ A.2 Technical Lemmas
The first attention layer finds all the edges in $G$. For adjacency list graph representation, we COPY the value of $n\times(i-1)$ (from the position $v_{i}$) to the positions $v_{i}^{1},\cdots,v_{i}^{m_{i}}$. Applying [10, Lemma C.7] and setting $\langle{\bm{q}}_{i},{\bm{k}}_{j}\rangle=\|\bm{x}_{i}-\bm{x}_{j}\|_{2}^{2}$, where $\bm{x}_{i}$ is the type of the node in the embedding suffices. For edge list graph representation, it suffices to COPY from the previous token. Thus we can set $\langle{\bm{q}}_{i},{\bm{k}}_{j}\rangle=(i-j-1)^{2}$.
@@
By [10, Lemma C.2], Equation 1 can be calculated with constant hidden dimension.
The second attention layer aggregates all $\mathsf{vec}(A(G[\{v_{i},v_{j}\}]))$ for each edge. We first calculate the MEAN of all valid $\mathsf{vec}(A(G[\{v_{i},v_{j}\}]))$ from the last layer by [10, Lemma C.8]. The result can be expressed as $\mathsf{vec}(A^{\prime}(G))$ where
@@
which can be implemented in the subsequent MLP by [10, Lemma C.2]. ∎
Lemma A.7 (One-Step Subgraph Isomorphism Indicator Tensor Calculation).
Fix integers $n,k,m\geq 1$. Let $V^{\prime}=\{v_{1}^{\prime},\dots,v_{k}^{\prime}\}$ be a set of $k$ vertices, and let ${\mathcal{F}}(V^{\prime})=(V_{1}^{\prime},V_{2}^{\prime},\dots,V_{m}^{\prime})$ be a $k$-node $m$-filtration on $V^{\prime}$ (defined in Definition 3.1). There exists $m$ two-layer MLPs with GeLU activation $\bm{f}_{1},\cdots,\bm{f}_{m}$ each with hidden dimension $O(n^{2}k^{2})$, such that for any directed graphs $G=(V,E)$ with $V=\{v_{1},\dots,v_{n}\}$ and $G^{\prime}=(V^{\prime},E^{\prime})$, $\bm{f}_{i}$ can output $\mathsf{vec}({\mathcal{T}}(G,G^{\prime}[V_{i}^{\prime}]))$ for input $\mathsf{vec}(A(G))\oplus\mathsf{vec}(A(G^{\prime}))\oplus\mathsf{vec}({\mathcal{T}}(G,G^{\prime}[V_{i-1}^{\prime}]))$ ($\oplus$ denotes concatenation), where $V_{0}^{\prime}:=\varnothing$.
@@
[…]
A.3 Proofs for Section 3
@@
Theorem A.8 (Formal Statement of Theorem 3.3).
Fix integers $n\geq k\geq 1,m\geq 1$. Let $V^{\prime}=\{v_{1}^{\prime},\dots,v_{k}^{\prime}\}$ be a set of $k$ vertices, and let ${\mathcal{F}}(V^{\prime})=(V_{1}^{\prime},V_{2}^{\prime},\dots,V_{m}^{\prime})$ be a $k$-node $m$-filtration on $V^{\prime}$. There exists a log-precision Transformer with $m+2$ layers, constant number of attention heads, and $O(n^{k})$ hidden dimension, such that for any directed graphs $G=(V,E)$ with $V=\{v_{1},\dots,v_{n}\}$ and $G^{\prime}=(V^{\prime},E^{\prime})$, the Transformer processing $G,G^{\prime}$ can output $\mathsf{vec}({\mathcal{T}}(G,G^{\prime}[V_{i}^{\prime}]))$ at layer $i+2$ for $i\in\{1,\cdots,m\}$. Here, $\mathsf{vec}(\cdot)$ is the vectorization for a tensor (formal defined in Definition A.3).
@@
The first two layers of Transformer calculates $\mathsf{vec}(A(G))$ and $\mathsf{vec}(A(G^{\prime}))$ for graph $G,G^{\prime}$. The desired output is at the last token for each graph, respectively. By Lemma A.6, this can be implemented with $O(n^{2})$ hidden dimension, regardless of the representation of $G,G^{\prime}$.
For the next $m$ layers, we first apply [10, Lemma C.7] to COPY $\mathsf{vec}(A(G)),\mathsf{vec}(A(G^{\prime}))$ in the attention layer. This can be implemented by adding special marks on the last token for each graph. In the subsequent MLP layer, we apply Lemma A.7 to calculate desired results with $O(n^{2}k^{2})$ hidden dimension. ∎
@@
Under 3.4, for directed graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$), a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension can output the unique $k$-tuple of vertices $(v_{i_{1}},\dots,v_{i_{k}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k}}=1$.
Theorem A.9 (Formal Statement of Theorem 3.5).
Fix integers $n\geq k\geq 1$, and let $V^{\prime}=\{v_{1}^{\prime},\dots,v_{k}^{\prime}\}$. There exists a log-precision Transformer with constant depth, constant number of attention heads, and $O(n^{k})$ hidden dimension, such that for any directed graphs $G=(V,E)$ with $V=\{v_{1},\dots,v_{n}\}$ and $G^{\prime}=(V^{\prime},E^{\prime})$ satisfying 3.4, the Transformer processing $G,G^{\prime}$ can output the unique tuple $(v_{i_{1}},\dots,v_{i_{k}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\cdots,i_{k}}=1$.
@@
We first take $m=1$ in Theorem 3.3, indicating that a constant depth Transformer can output $\mathsf{vec}({\mathcal{T}}(G,G^{\prime}))$ for any directed graphs $G,G^{\prime}$. By 3.4, $\text{ReLU}(\mathsf{vec}({\mathcal{T}}(G,G^{\prime})))$ is a one-hot vector and can be obtained with a two-layer MLP via [10, Lemma C.2].
@@
Theorem 3.7 (Expressiveness for Multi-Num Extraction).
For directed graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$), a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension can output all $k$-tuples of vertices $(v_{i_{1}},\dots,v_{i_{k}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k}}=1$.
Theorem A.10 (Formal Statement of Theorem 3.7).
@@
Follow the proof of Theorem 3.5, we first calculate $\bm{v}^{1}=\text{ReLU}(\mathsf{vec}({\mathcal{T}}(G,G^{\prime})))$ which marks all feasible tuples with $1$, and 00 otherwise.
Next, we calculate $\bm{v}^{2}$ where $\bm{v}^{2}_{i}=\bm{v}^{1}_{1}+\cdots+\bm{v}^{1}_{i}$ by linear projection, and define $\bm{v}^{3}_{i}=\bm{v}^{1}_{i}\bm{v}_{i}^{2}$. In $\bm{v}^{3}_{i}$, the feasible tuples are marked as $1,2,\cdots$, and 00 otherwise. By [10, Lemma C.1], $\bm{v}^{2},\bm{v}^{3}$ can be obtained with a two-layer MLP.
@@
then following with the proofs in Theorem 3.5 to output the current position. By [10, Lemma C.2], the above steps can be obtained with constant-layer MLPs, which concludes our proof. ∎
Theorem 3.8 (Expressiveness for Multi-Shape Extraction).
For a directed graph $G=(V,E)$ ($|V|=n$) and any target subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with $|V^{\prime}|=k^{\prime}\leq k$ satisfying 3.4, a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension can output the unique $k^{\prime}$-tuple of vertices $(v_{i_{1}},\dots,v_{i_{k^{\prime}}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k^{\prime}}}=1$.
Theorem A.11 (Formal Statement of Theorem 3.8).
Fix integers $n\geq k\geq 1$. There exists a log-precision Transformer with constant depth, constant number of attention heads, and $O(n^{k})$ hidden dimension, such that for any directed graph $G=(V,E)$ (with $V=\{v_{1},\dots,v_{n}\}$) and $G^{\prime}=(V^{\prime},E^{\prime})$ (with $V^{\prime}=\{v_{1}^{\prime},\cdots,v_{k^{\prime}}^{\prime}\}$ where $k^{\prime}\leq k$) satisfying 3.4, the Transformer processing $G,G^{\prime}$ can output the unique tuples $(v_{i_{1}},\dots,v_{i_{k^{\prime}}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\cdots,i_{k^{\prime}}}=1$.

### v2

@@ 3.2.2 Question prompt encoders
4 Consistency in LLM Graph Understanding Behavior
As discussed in Section˜3, we identify three key findings: In terms of the Transformer mechanism, (1) Transformers perform the ISF process to extract substructures simultaneously. Regarding input formulation, (2) both EL and AL can represent the adjacency matrix $A(G)$, though EL may perform slightly worse than AL due to sequence length limitations, and (3) Transformers tend to abstract substructures into a sequence of tokens in the question prompt, rather than fully capturing the underlying topological concepts. Since decoder-only Transformers are a common architecture in modern LLMs, we further evaluate whether LLMs exhibit the ISF process during substructure extraction, and whether our understanding of input formulation can explain their behavior.
@@ 4 Consistency in LLM Graph Understanding Behavior
To investigate whether LLMs exhibit the ISF process, we visualize the fine-tuned LLaMA 3.1-8B-Instruct model on a triangle detection task (details in Section˜E.4.2), as shown in Figure˜11. The results align with our findings in Section˜3.1: the model often identifies the correct answer before generating it, and deeper layers better distinguish between similar answers. For example, the responses of the substructures in node ID ‘302’ and ‘304’ become more separable at layer 23 than at layer 17. We quantify this trend in Figure˜11, where Adjusted Rand Index (ARI) and Normalized Mutual Information (NMI) scores increase with depth. However, unlike Transformers trained from scratch, which only predict answers, LLMs can generate explanatory content. For instance, 23% of responses include Python code, while others describe node relationships in the graph. As a result, ARI and NMI scores slightly decrease in the ending layers.
@@
5 Understanding on Complex Graphs
In Section˜3, we highlight the ISF process as a key mechanism enabling Transformers to solve the substructure extraction task. Moreover, we observe that when text-based descriptions can be mapped to the adjacency matrix $A(G)$, Transformers can perform substructure extraction on the given graph. Building on these insights, we explore broader applications: in Section˜5.1, we introduce a new method for efficiently reasoning over composite substructures and in Section˜5.2, we examine how Transformers adapt to attributed graphs.
@@ 5.1 Thinking in substructures
As introduced in Section 3, Transformers apply the ISF process to extract substructures. This process runs synchronously across different patterns, with smaller ones being easier to detect. Besides, complex structures often consist of simpler components, which we refer to as decomposing substructures. Building on these observations, we propose the Thinking-in-Substructure (Tins) method to explain how decoder-only Transformers solve complex substructure reasoning tasks. For example, [7] reports that reasoning language models decompose a house pattern into a triangle and a square to search for the substructures. We reformulate the answer generation part as $)\mathsf{ANS_{Tins}}(\hat{G})=(\{P_{1}\},\{P_{2}\},\dots,\{P_{t}\},\mathrm{<ANS>},\mathsf{ANS}(\hat{G}$)), where $\{P_{i}\}$ is the collection of each decomposing structure and $\mathrm{<ANS>}$ is a special token to indicate that the followings are final answers. This decomposition reduces the extraction complexity from $O(n^{k})$ to $O(n^{q})$, where $q$ and $k$ are the maximum size of decomposing substructures and target substructures, respectively, with $q<k$. We show the proof in Theorem˜D.13 in Section˜D.4.
To verify the efficiency of Tins, in the experiment, we design 4 different composite substructures with their decomposition process trained with 100K samples. The experiment settings are in Section˜E.4.3. The results suggest that Tins can help transformers significantly improve the performance with limited training data. The overall performance can increase 10%, and in the 3 layer for diagonal structure, the performance increases about 46% percent.
@@
Substructures Directly Preds Decomposition Tins
4 layer 3 layer 4 layer 3 layer
Diagonal [Uncaptioned image] 0.6314 0.1998 [Uncaptioned image] 0.8648 0.6606
@@ 5.2 Attributed graphs
As shown in Theorem˜3.5, the only condition required for Transformers to extract substructures is $\mathcal{T}(G,G^{\prime})=1$. Therefore, if node features are uniquely assigned to ensure a distinct graph representation for each graph, Transformers can extract substructures while incorporating these features, as discussed in Theorem˜D.16. Taking the AL-based feature description as an example, we define the attributed graph representation as $\mathsf{AL_{f}}(G)=(v_{1}f_{1};\text{``:''};v_{1}^{1}f_{1}^{1};\cdots;v_{1}^{m_{1}}f_{1}^{m_{1}};\text{``,''};\cdots;\text{``,''};v_{n}f_{n};\text{``:''};v_{n}^{1}f_{n}^{1};\cdots;v_{n}^{m_{n}}f_{n}^{m_{n}}).$ , where $f_{i}$ is the node features.
@@ 6 Related work
The mechanisms originate from graph learning methods, with detailed comparisons of graph neural networks, graph transformers, and decoder-only transformers in Appendix˜C. Theory now probes how vanilla Transformers perform graph reasoning. Log-depth models suffice, and are necessary, for connectivity and cycles [17], while width can trade for depth [29]. ALPINE [24] shows a GPT layer embeds adjacency and reachability, validating on planning tasks; two-layer decoders trained on shortest-path learn spectral line-graph embeddings instead of Dijkstra-style rules [5]. Surveys relate attention power to Weisfeiler–Lehman bounds and over-squashing limits [13, 19]; scalable variants such as AnchorGT mitigate $O(n^{2})$ cost without losing accuracy [34]. Hierarchical distances or sparse global attention keep Transformers competitive on large or molecular graphs [9, 8, 30]. Collectively, these studies view attention heads as induced-neighbourhood selectors—exactly the mechanism ISF formalises via filtration depth.
@@ 7 Conclusion
Acknowledgement
Xinnan Dai, Jay Revolinsky, Kai Guo, and Jiliang Tang are supported by the National Science Foundation (NSF) under grant numbers CNS2321416, IIS2212032, IIS2212144, IIS 2504089, DUE2234015, CNS2246050, DRL2405483 and IOS2035472, the Michigan Department of Agriculture and Rural Development, US Dept of Commerce, Gates Foundation, Amazon Faculty Award, Meta, NVIDIA, Microsoft and SNAP.
Appendix A Boarder Impact
In this work, we present new insights into how Transformers solve the substructure extraction task, offering a deeper understanding of their internal mechanisms. Our contributions span three key areas: (1) we introduce a novel concept of Induced Substructure Filtration (ISF), instructing LLMs in structure data understanding; (2) we propose a decomposition-based approach for tackling complex substructure reasoning by breaking down intricate patterns into simpler components, enabling more efficient extraction. This concept of thinking in substructures can generalize beyond graph tasks—for example, complementing step-by-step reasoning with pattern-by-pattern thinking; and (3) we provide both theoretical and empirical evidence supporting the development of graph foundation models, highlighting the potential of Transformers as backbones for structured learning tasks.
Appendix B Limitation
In this work, we focus primarily on the fundamentals of the substructure extraction task. However, other substructure-related tasks remain to be explored in future research. Additionally, our current study provides a high-level overview of decoder-only Transformers, leaving out theoretical details. Future work can extend this foundation to develop a more comprehensive and rigorous understanding.
Appendix C Related work
The mechanisms by which machine learning models learn graph-related problems have been widely studied, ranging from graph neural networks to transformers. However, due to their differing input–output formulations, the underlying mechanisms vary substantially, as summarized in Table˜5.
Table 5: Comparison among GNNs, Graph Transformers, and Decoder-Only Transformers (LLMs).
GNNs Graph Transformer Decoder-Only Transformer
Input Discrete graph data (node features, adjacency matrix) Discrete graph data (node features, adjacency matrix) Textual description of a graph
Output Scalar value (e.g., count) or fixed-size vector (e.g., for classification) Scalar value (e.g., count) or fixed-size vector (e.g., for classification) Text tokens forming a human-readable structural description in indefinite length
Learning Formulation Encode graph via message passing and neighborhood aggregation; supervised on scalar outputs Encode graph via graph-aware self-attention; supervised on scalar outputs Next-token prediction over graph-structured text
Mechanism 1-WL [4] k-WL [14] / SEG-WL [33] The proposed ISF (Ours)
Appendix D More on Theoretical Analysis
D.1 Preliminaries
@@
Definition D.1 (Adjacency List Graph Representation).
@@
Definition D.2 (Edge List Graph Representation).
@@
Definition D.3 (Tensor Vectorization).
@@
Definition D.5 (Subgraph Isomorphism Indicator Tensor).
@@
[…]
@@
Throughout our theoretical analysis, we consider log-precision auto-regressive Transformer, instead of constant-precision Transformer. See [12, Appendix B] for more discussions.
D.2 Technical Lemmas
Lemma D.6 (Adjacency Matrix Extraction).
@@ D.2 Technical Lemmas
The first attention layer finds all the edges in $G$. For adjacency list graph representation, we COPY the value of $n\times(i-1)$ (from the position $v_{i}$) to the positions $v_{i}^{1},\cdots,v_{i}^{m_{i}}$. Applying [11, Lemma C.7] and setting $\langle{\bm{q}}_{i},{\bm{k}}_{j}\rangle=\|\bm{x}_{i}-\bm{x}_{j}\|_{2}^{2}$, where $\bm{x}_{i}$ is the type of the node in the embedding suffices. For edge list graph representation, it suffices to COPY from the previous token. Thus we can set $\langle{\bm{q}}_{i},{\bm{k}}_{j}\rangle=(i-j-1)^{2}$.
@@
By [11, Lemma C.2], Equation˜1 can be calculated with constant hidden dimension.
The second attention layer aggregates all $\mathsf{vec}(A(G[\{v_{i},v_{j}\}]))$ for each edge. We first calculate the MEAN of all valid $\mathsf{vec}(A(G[\{v_{i},v_{j}\}]))$ from the last layer by [11, Lemma C.8]. The result can be expressed as $\mathsf{vec}(A^{\prime}(G))$ where
@@
which can be implemented in the subsequent MLP by [11, Lemma C.2]. ∎
Lemma D.7 (One-Step Subgraph Isomorphism Indicator Tensor Calculation).
Fix integers $n,k,m\geq 1$. Let $V^{\prime}=\{v_{1}^{\prime},\dots,v_{k}^{\prime}\}$ be a set of $k$ vertices, and let ${\mathcal{F}}(V^{\prime})=(V_{1}^{\prime},V_{2}^{\prime},\dots,V_{m}^{\prime})$ be a $k$-node $m$-filtration on $V^{\prime}$ (defined in Definition˜3.1). There exists $m$ two-layer MLPs with GeLU activation $\bm{f}_{1},\cdots,\bm{f}_{m}$ each with hidden dimension $O(n^{2}k^{2})$, such that for any directed graphs $G=(V,E)$ with $V=\{v_{1},\dots,v_{n}\}$ and $G^{\prime}=(V^{\prime},E^{\prime})$, $\bm{f}_{i}$ can output $\mathsf{vec}({\mathcal{T}}(G,G^{\prime}[V_{i}^{\prime}]))$ for input $\mathsf{vec}(A(G))\oplus\mathsf{vec}(A(G^{\prime}))\oplus\mathsf{vec}({\mathcal{T}}(G,G^{\prime}[V_{i-1}^{\prime}]))$ ($\oplus$ denotes concatenation), where $V_{0}^{\prime}:=\varnothing$.
@@
Thus it suffices to calculate the value of $A(G^{\prime})_{x,y}A(G)_{x^{\prime},y^{\prime}}$ for $1\leq x,y\leq k,\;1\leq x^{\prime},y^{\prime}\leq n$. By [11, Lemma C.1], this can be implemented with $O(n^{2}k^{2})$ hidden dimension. ∎
D.3 Proofs for Section˜3
@@
Theorem D.8 (Formal Statement of Theorem˜3.3).
Fix integers $n\geq k\geq 1,m\geq 1$. Let $V^{\prime}=\{v_{1}^{\prime},\dots,v_{k}^{\prime}\}$ be a set of $k$ vertices, and let ${\mathcal{F}}(V^{\prime})=(V_{1}^{\prime},V_{2}^{\prime},\dots,V_{m}^{\prime})$ be a $k$-node $m$-filtration on $V^{\prime}$. There exists a log-precision Transformer with $m+2$ layers, constant number of attention heads, and $O(n^{k})$ hidden dimension, such that for any directed graphs $G=(V,E)$ with $V=\{v_{1},\dots,v_{n}\}$ and $G^{\prime}=(V^{\prime},E^{\prime})$, the Transformer processing $G,G^{\prime}$ can output $\mathsf{vec}({\mathcal{T}}(G,G^{\prime}[V_{i}^{\prime}]))$ at layer $i+2$ for $i\in\{1,\cdots,m\}$. Here, $\mathsf{vec}(\cdot)$ is the vectorization for a tensor (formal defined in Definition˜D.3).
@@
The first two layers of Transformer calculates $\mathsf{vec}(A(G))$ and $\mathsf{vec}(A(G^{\prime}))$ for graph $G,G^{\prime}$. The desired output is at the last token for each graph, respectively. By Lemma˜D.6, this can be implemented with $O(n^{2})$ hidden dimension, regardless of the representation of $G,G^{\prime}$.
For the next $m$ layers, we first apply [11, Lemma C.7] to COPY $\mathsf{vec}(A(G)),\mathsf{vec}(A(G^{\prime}))$ in the attention layer. This can be implemented by adding special marks on the last token for each graph. In the subsequent MLP layer, we apply Lemma˜D.7 to calculate desired results with $O(n^{2}k^{2})$ hidden dimension. ∎
@@
Under ˜3.4, for directed graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$), a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension can output the unique $k$-tuple of vertices $(v_{i_{1}},\dots,v_{i_{k}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\dots,i_{k}}=1$.
Theorem D.9 (Formal Statement of Theorem˜3.5).
Fix integers $n\geq k\geq 1$, and let $V^{\prime}=\{v_{1}^{\prime},\dots,v_{k}^{\prime}\}$. There exists a log-precision Transformer with constant depth, constant number of attention heads, and $O(n^{k})$ hidden dimension, such that for any directed graphs $G=(V,E)$ with $V=\{v_{1},\dots,v_{n}\}$ and $G^{\prime}=(V^{\prime},E^{\prime})$ satisfying ˜3.4, the Transformer processing $G,G^{\prime}$ can output the unique tuple $(v_{i_{1}},\dots,v_{i_{k}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\cdots,i_{k}}=1$.
@@
We first take $m=1$ in Theorem˜3.3, indicating that a constant depth Transformer can output $\mathsf{vec}({\mathcal{T}}(G,G^{\prime}))$ for any directed graphs $G,G^{\prime}$. By ˜3.4, $\text{ReLU}(\mathsf{vec}({\mathcal{T}}(G,G^{\prime})))$ is a one-hot vector and can be obtained with a two-layer MLP via [11, Lemma C.2].
@@
Theorem 3.8 (Expressiveness for Single-Shape-Multi-Num Extraction).
Fix integers $n\geq k\geq 1$. There exists a log-precision Transformer with constant depth, constant number of attention heads, and $O(n^{k})$ hidden dimension that can complete Single-Shape-Multi-Num Extraction defined in Definition˜3.7 for directed graphs $G=(V,E)$ ($|V|=n$) and $G^{\prime}=(V^{\prime},E^{\prime})$ ($|V^{\prime}|=k$).
Theorem D.10 (Formal Statement of Theorem˜3.8).
@@
Follow the proof of Theorem˜3.5, we first calculate $\bm{v}^{1}=\text{ReLU}(\mathsf{vec}({\mathcal{T}}(G,G^{\prime})))$ which marks all feasible tuples with $1$, and 0 otherwise.
Next, we calculate $\bm{v}^{2}$ where $\bm{v}^{2}_{i}=\bm{v}^{1}_{1}+\cdots+\bm{v}^{1}_{i}$ by linear projection, and define $\bm{v}^{3}_{i}=\bm{v}^{1}_{i}\bm{v}_{i}^{2}$. In $\bm{v}^{3}_{i}$, the feasible tuples are marked as $1,2,\cdots$, and 0 otherwise. By [11, Lemma C.1], $\bm{v}^{2},\bm{v}^{3}$ can be obtained with a two-layer MLP.
@@
then following with the proofs in Theorem˜3.5 to output the current position. By [11, Lemma C.2], the above steps can be obtained with constant-layer MLPs, which concludes our proof. ∎
Theorem 3.10 (Expressiveness for Multi-Shape-Single-Num Extraction).
Fix integers $n\geq k\geq 1$. There exists a log-precision Transformer with constant depth, constant heads, and $O(n^{k})$ hidden dimension that can complete Multi-Shape-Single-Num Extraction defined in Definition˜3.9 for a directed graph $G=(V,E)$ ($|V|=n$) and any target subgraph $G^{\prime}=(V^{\prime},E^{\prime})$ with $|V^{\prime}|=k^{\prime}\leq k$ satisfying ˜3.4.
Theorem D.11 (Formal Statement of Theorem˜3.10).
Fix integers $n\geq k\geq 1$. There exists a log-precision Transformer with constant depth, constant number of attention heads, and $O(n^{k})$ hidden dimension, such that for any directed graph $G=(V,E)$ (with $V=\{v_{1},\dots,v_{n}\}$) and $G^{\prime}=(V^{\prime},E^{\prime})$ (with $V^{\prime}=\{v_{1}^{\prime},\cdots,v_{k^{\prime}}^{\prime}\}$ where $k^{\prime}\leq k$) satisfying ˜3.4, the Transformer processing $G,G^{\prime}$ can output the unique tuples $(v_{i_{1}},\dots,v_{i_{k^{\prime}}})$ for which ${\mathcal{T}}(G,G^{\prime})_{i_{1},\cdots,i_{k^{\prime}}}=1$.